# Leveraging Causal Reasoning Method for Explaining Medical Image Segmentation Models

### v1

@@ 3.1 Preliminaries
We model interventions on images within the framework of the potential outcome model [37]. Ideally, the difference between potential outcomes can be interpreted as the causal effect of a treatment on the result. For instance, given a binary treatment $T=0/1$, each individual subject $I$ can receive only one treatment in practice, denoted by the causal intervention $do(I=0/1)$ [38], leading to corresponding outcomes $R_{0/1}$. The individual treatment effect (ITE) is then defined as $R_{1}-R_{0}$. Due to individual variability or other confounding factors, this estimate may be biased. To mitigate such effects, the analysis can be extended to a population of size $N$, allowing for the computation of the ATE as a more robust measure:
@@
Here, $I$ corresponds to the image $X$, and $R$ corresponds to the evaluation metric $M$ in the visual task formulation.
@@ 3.3 Principles of Intervention
To mitigate limitations, we opt to extract perturbations from the natural distribution corresponding to the dataset. Specifically, when perturbing a dermatological image, we first partition a subset from the dataset that excludes the target image. The intervention $do(\cdot)$ is then defined by randomly cropping and extracting a block $b$ from this subset to replace the original region. In the segmentation setting, the output $Y$ must be evaluated by the model $M$. As illustrated in Fig. 1(a) and (b), let the model’s response to the initial RoI be denoted as $M_{0}={M_{RoI}}({f_{\theta}}(X))$. The perturbation applied to the $i$-th patch is represented as $\left.X\right|do({X_{p_{i}}}=b)$. ${\rm{ITE}}{p_{i}}$ can be computed and substituted into Eq. 1. By performing $N$ such interventions, ${\rm{ATE}}{p_{i}}$ is computed as:
@@
where $K$ denotes the number of patches the image is divided into, $N$ is a user-defined parameter, and $M$ can be instantiated with any commonly used segmentation metric. In this work, we adopt the standard Dice similarity coefficient (DSC) [41].
@@ 3.4.1 Coarse-to-Fine Patch Screening
The causal effect of the RoI on itself is infinite, but it is not sensitive to perturbations at every other location. In particular, changes in pixels that are far away may have little influence on the model’s judgment of the RoI. Therefore, interventions can be performed in a coarse-to-fine manner. Specifically, a patch is first substituted $S$ times (set to $3$ in this work). If all the resulting ITE values are below the threshold $\tau$ (set to $0.02$ in this work), the patch is considered irrelevant to the RoI and thus ignored. According to Fig. 1(c), all patches deemed relevant will proceed to the full ATE calculation:
@@ 3.4.2 Convergence Steps
A highly confident causal relationship may require many inference steps, which inevitably prolongs the overall procedure. Therefore, it is necessary to determine a value of $N$ beyond which further increases yield diminishing returns and do not introduce significant bias. Ultimately, $N=50$ was chosen, as shown in Fig. 3. The inference results initially fluctuate dramatically but gradually stabilize as the number of steps increases. It can be observed that around step 50, the DSC ATE for the majority of networks converges, indicating that additional steps beyond this point do not lead to substantial differences.

### v2

@@ 3.1 Preliminaries
We model interventions on images within the framework of the potential outcome model [37]. Ideally, the difference between potential outcomes can be interpreted as the causal effect of a treatment on the result. For instance, given a binary treatment $T=0/1$, each individual subject II can receive only one treatment in practice, denoted by the causal intervention $do(I=0/1)$ [38], leading to corresponding outcomes $R_{0/1}$. The individual treatment effect (ITE) is then defined as $R_{1}-R_{0}$. Due to individual variability or other confounding factors, this estimate may be biased. To mitigate such effects, the analysis can be extended to a population of size NN, allowing for the computation of the ATE as a more robust measure:
@@
Here, II corresponds to the image $X$, and $R$ corresponds to the evaluation metric $M$ in the visual task formulation.
@@ 3.3 Principles of Intervention
To mitigate limitations, we opt to extract perturbations from the natural distribution corresponding to the dataset. Specifically, when perturbing a dermatological image, we first partition a subset from the dataset that excludes the target image. The intervention $do(\cdot)$ is then defined by randomly cropping and extracting a block $b$ from this subset to replace the original region. In the segmentation setting, the output $Y$ must be evaluated by the model $M$. As illustrated in Fig. 1(a) and (b), let the model’s response to the initial RoI be denoted as $M_{0}={M_{RoI}}({f_{\theta}}(X))$. The perturbation applied to the $i$-th patch is represented as $\left.X\right|do({X_{p_{i}}}=b)$. ${\rm{ITE}}{p_{i}}$ can be computed and substituted into Eq. 1. By performing NN such interventions, ${\rm{ATE}}{p_{i}}$ is computed as:
@@
where $K$ denotes the number of patches the image is divided into, NN is a user-defined parameter, and $M$ can be instantiated with any commonly used segmentation metric. In this work, we adopt the standard Dice similarity coefficient (DSC) [41].
@@ 3.4.1 Coarse-to-Fine Patch Screening
The causal effect of the RoI on itself is infinite, but it is not sensitive to perturbations at every other location. In particular, changes in pixels that are far away may have little influence on the model’s judgment of the RoI. Therefore, interventions can be performed in a coarse-to-fine manner. Specifically, a patch is first substituted $S$ times (set to 33 in this work). If all the resulting ITE values are below the threshold $\tau$ (set to $0.02$ in this work), the patch is considered irrelevant to the RoI and thus ignored. According to Fig. 1(c), all patches deemed relevant will proceed to the full ATE calculation:
@@ 3.4.2 Convergence Steps
A highly confident causal relationship may require many inference steps, which inevitably prolongs the overall procedure. Therefore, it is necessary to determine a value of NN beyond which further increases yield diminishing returns and do not introduce significant bias. Ultimately, $N=50$ was chosen, as shown in Fig. 3. The inference results initially fluctuate dramatically but gradually stabilize as the number of steps increases. It can be observed that around step 50, the DSC ATE for the majority of networks converges, indicating that additional steps beyond this point do not lead to substantial differences.